# Measurement of ion acceleration and diffusion in a laser-driven magnetized plasma

### v1

@@ Abstract
Here we present results from an experiment performed at the GSI Helmholtz Centre for Heavy Ion Research. A mono-energetic beam of chromium ions with initial energies of MeV was fired through a magnetized interaction region formed by the collision of two counter-propagating laser-ablated plasma jets. While laser interferometry revealed the absence of strong fluid-scale turbulence, acceleration and diffusion of the beam ions was driven by wave-particle interactions. A possible mechanism is particle acceleration by electrostatic, short scale length kinetic turbulence, such as the lower-hybrid drift instability.
@@
Experimental setup— The experimental platform employed for the experiment at the GSI Helmholtz Institute for Heavy Ion research is shown schematically in Fig. 1. Supersonic plasma jets are produced via laser ablation of two opposing target foils, each machined with an embedded grid structure to impose density perturbations. Simultaneously, seed magnetic fields in the azimuthal direction are generated by the Biermann battery mechanism from the cross-product of the density and temperature gradients [25, 26, 27]. The two jets collide at the midpoint between the two foils, defined as the target chamber center (TCC), forming an approximately cylindrical magnetized interaction region. The intention of the design was for the initial density perturbations advected by the jets to interpenetrate, introducing shearing motions which would drive turbulent mixing of the plasma. The scale of these instabilities is however, predicted to be smaller than the experimental resolution of the diagnostics [24]. Additionally, by only driving a single target foil, the case of reduced fluctuations, plasma density, and field strength was accessed for comparison. Interferometry and ion deflectrometry were fielded as diagnostics for the experimental plasma conditions.
The monoenergetic ion beam from the UNIversal Linear ACcelerator (UNILAC) was directed through the interaction region, acting as a surrogate for CR particles. The UNILAC supplied a Gaussian pulse train (4 ns width, 28 ns spacing), with only one pulse traversing the interaction region in each shot. The energy and initial charge state of the ion beam were tunable parameters and we present results obtained using the following: 52Cr14+ with energies of 8.75 MeV/u; 52Cr20+ with energies of 8.661 MeV/u.
@@
Results— Line-integrated electron densities in the ablated plasma flows and interaction region were measured using laser interferometry, providing a coarse measurement of the plasma density structure. This data also provided an upper bound for the level of fluid turbulence by considering finer density fluctuations below the smaller scale width of the plasma ( mm). In addition, the density profile of the plasma expansion was used to estimate the temperature.
@@
The free electron density, , is retrieved from the raw interferograms via the following process. Firstly, the fringes were traced manually to form phase contour maps. In regions associated with strong density gradients, for example adjacent to the target foils, individual fringes cannot be resolved – these regions are excluded from the interpolation.
@@
The evolution of the inflow plasma jets was captured by interferometry on shots with only one target foil driven by nhelix. The expanding jet can be approximated as isothermal, since the ratio of the temperature equilibration timescale to the experiment duration can be shown to be [32] ; equivalently this implies a small Péclet number (). An isothermal expansion has a known density profile [33, 34] dependent on the plasma sound speed . Since is related to the plasma temperature [35], at a time of ns, approximately the time the jets collide, the temperature of the inflow jets from this analysis is eV. Even accounting for collisional heating, we can expect the interaction region plasma to be in the temperature regime of the low hundreds of eV.
@@
Spatially resolved ion fluence was recorded with this setup. The distribution of pits demarcating the position of ion impacts consists of a saturated central region containing the bulk of the ion beam and a surrounding ‘halo’ decreasing in pit density with increasing radial distance; Fig. 4A and 4B show a section of the data aligned radially outwards from a knife edge positioned mm from the center of the ion beam, for shots without and with plasma respectively. As the ion beam traverses the plasma, it accumulates a random-walk deflection proportional to the magnetic field strength, which dominates over contributions from the electric fields and Coulomb collisions. The relative contributions from magnetic and electric fields was compared by considering the Lorentz equation of motion and applying resistive MHD; the ratio of deflection force from electric and magnetic fields , where is the characteristic fluid velocity and the initial ion beam velocity.
@@
where , and are the charge state, mass and energy of the beam ions. An upper estimate for the magnetic coherence length is found by taking the grid spacing (60 um) on the target foils and accounting for the approximate doubling in spatial extent of the plasma: mm. This is on the order of the outer scale of turbulence, which in previous experiments (on similar setups) has been larger than [19]. Using the experimental value of obtained from the width of the distribution in Fig. 4D, this gives a minimum field strength of kG. An upper bound for the field strength can be defined using a simple geometric argument: assuming all the counts in the distribution of Fig. 4D come from ions at the center of the beam, we obtain which is found by summing the initial ion beam divergence and the measured . This corresponds to a field strength of kG. Thus, we estimate kG, which has implications for later discussion on acceleration from fluid and short-scale electrostatic turbulence.
@@
where is the plasma frequency and is the Coulomb logarithm. Using the experimentally constrained plasma conditions in Table I, the electron-ion stopping dominates, with values of and MeV for and , respectively, assuming a path length of 2 mm. This magnitude of deceleration cannot account for the observed shifts in mean energy. Similarly, the expected parallel diffusion associated with Coulomb collisions is found to be negligible ( eV) under the experimental conditions, due to the high initial velocity of the ion beam.
@@
where is the speed of light and and are the RMS velocity and magnetic field, respectively. Taking (bounded by deflectrometry) and (bounded by interferometry), the maximum associated with second-order Fermi acceleration (see Table I) is much smaller than the initial energy spread of the ion beam ( MeV) and the energy resolution of the diamond detector. Consequently, while this mechanism may contribute, it cannot be responsible for the changes measured.
@@
The observation of bulk energisation cannot be explained by the broadening process outlined above alone, and suggests that coherent fields which facilitate particle acceleration are generated in the plasma. A tantalising possibility consistent with our experiment is the role of short scale electrostatic turbulence (below the interferometry diagnostic’s resolution) in developing coherent structures within the plasma. In literature, both Langmuir wave [48] and LHW [49, 50] turbulence have been shown to drive the evolution of these wave structures.

### v2

@@ Abstract
Abstract— Here we present results from an experiment performed at the GSI Helmholtz Centre for Heavy Ion Research. A mono-energetic beam of chromium ions with initial energies of MeV was fired through a magnetized interaction region formed by the collision of two counter-propagating laser-ablated plasma jets. While laser interferometry revealed the absence of strong fluid-scale turbulence, acceleration and diffusion of the beam ions was driven by wave-particle interactions. A possible mechanism is particle acceleration by electrostatic, short scale length kinetic turbulence, such as the lower-hybrid drift instability.
Introduction
@@
Results
The experimental platform employed for the experiment at the GSI Helmholtz Institute for Heavy Ion research is shown schematically in Fig. 1. Supersonic plasma jets are produced via laser ablation of two opposing target foils, each machined with an embedded grid structure to impose density perturbations. Simultaneously, seed magnetic fields in the azimuthal direction are generated by the Biermann battery mechanism from the cross-product of the density and temperature gradients [25, 26, 27]. The two jets collide at the midpoint between the two foils, defined as the target chamber center (TCC), forming an approximately cylindrical magnetized interaction region. The intention of the design was for the initial density perturbations advected by the jets to interpenetrate, introducing shearing motions which would drive turbulent mixing of the plasma. The scale of these instabilities is however, predicted to be smaller than the experimental resolution of the diagnostics [24]. Additionally, by only driving a single target foil, the case of reduced fluctuations, plasma density, and field strength was accessed for comparison. Interferometry and ion deflectrometry were fielded as diagnostics for the experimental plasma conditions.
The monoenergetic ion beam from the UNIversal Linear ACcelerator (UNILAC) was directed through the interaction region, acting as a surrogate for CR particles. The UNILAC supplied a Gaussian pulse train (4 ns width, 28 ns spacing), with only one pulse traversing the interaction region in each shot. Each pulse contains approximately ions, a sufficiently low number such that it can be treated as a test beam, with no collective effects expected. The energy and initial charge state of the ion beam were tunable parameters and we present results obtained using the following: 52Cr14+ with energies of 8.75 MeV/u; 52Cr20+ with energies of 8.661 MeV/u.
@@ Results
Line-integrated electron densities in the ablated plasma flows and interaction region were measured using laser interferometry, providing a coarse measurement of the plasma density structure. This data also provided an upper bound for the level of fluid turbulence by considering finer density fluctuations below the smaller scale width of the plasma ( mm). In addition, the density profile of the plasma expansion was used to place a conservative lower bound on the temperature.
@@
The free electron density, , is retrieved from the raw interferograms via the following process. Firstly, the fringes were traced manually to form phase contour maps (Fig. 2A and B). In regions associated with strong density gradients, for example adjacent to the target foils, individual fringes cannot be resolved – these regions are excluded from the interpolation.
@@
The evolution of the inflow plasma jets was captured by interferometry on shots with only one target foil driven by nhelix, as seen in Fig. 3. The expanding jet can be approximated as isothermal, since the ratio of the temperature equilibration timescale to the experiment duration can be shown to be [32] ; equivalently this implies a small Péclet number (). An isothermal expansion has a known density profile [33, 34] dependent on the plasma sound speed . Since is related to the plasma temperature [35], at a time of ns, approximately the time the jets collide, the temperature of the inflow jets from this analysis is eV. Even accounting for collisional heating, we can expect the interaction region plasma to be in the temperature regime of the low hundreds of eV.
@@
Spatially resolved ion fluence was recorded with this setup. The distribution of pits demarcating the position of ion impacts consists of a saturated central region containing the bulk of the ion beam and a surrounding ‘halo’ decreasing in pit density with increasing radial distance; Fig. 4A and 4B show a section of the data aligned radially outwards from a knife edge positioned mm from the center of the ion beam, for shots without and with plasma respectively. As the ion beam traverses the plasma, it accumulates a random-walk deflection proportional to the magnetic field strength, which dominates over contributions from the electric fields and Coulomb collisions. The relative contributions from magnetic and electric fields was compared by considering the Lorentz equation of motion and applying resistive MHD; the ratio of deflection force from electric and magnetic fields , where is the characteristic fluid velocity () bounded by estimating the plasma jet displacement between subsequent interferograms, and the initial ion beam velocity.
@@
where , and are the charge state, mass and energy of the beam ions. The magnetic coherence length is the characteristic length over which the stochastic magnetic fields maintain coherency and is essentially a measure of the size of the stochastic field structures. An upper estimate for is found by taking the grid spacing (60 µm) on the target foils and accounting for the approximate doubling in spatial extent of the plasma: mm. This is on the order of the outer scale of turbulence, which in previous experiments (on similar setups) has been larger than [19]. Using the experimental value of obtained from the width of the distribution in Fig. 4D, this gives a minimum field strength of kG. An upper bound for the field strength can be defined using a simple geometric argument: assuming all the counts in the distribution of Fig. 4D come from ions at the center of the beam, we obtain which is found by summing the initial ion beam divergence and the measured . This corresponds to a field strength of kG. Thus, we estimate kG, which has implications for later discussion on acceleration from fluid and short-scale electrostatic turbulence.
@@
where is the initial ion beam velocity, is the plasma frequency, and is the Coulomb logarithm. Using the experimentally constrained plasma conditions in Table I, the electron-ion stopping dominates, with values of and MeV for and , respectively, assuming a path length of 2 mm. This magnitude of deceleration cannot account for the observed shifts in mean energy. Similarly, the expected parallel diffusion associated with Coulomb collisions is found to be negligible ( eV) under the experimental conditions, due to the high initial velocity of the ion beam.
@@
where is the speed of light and and are the RMS velocity and magnetic field, respectively. Taking (bounded by deflectrometry) and (bounded by interferometry), the maximum associated with second-order Fermi acceleration (see Table I) is much smaller than the initial energy spread of the ion beam ( MeV) and the energy resolution of the diamond detector. As and hence are only bounded from below by interferometry, one might consider the effect of higher temperature on the magnitude of . A conservative back-of-the-envelope scaling () shows that one would need temperatures of order before the Fermi term could become competitive – clearly not physical for this platform and inconsistent with the observed density and expansion dynamics. Consequently, while this mechanism may contribute, it cannot be responsible for the changes measured – a conclusion that is robust to substantially higher local temperatures.
@@
An alternative avenue for the generation of electrostatic waves near the lower-hybrid frequency is via beam-driven instabilities, such as the modified two-stream instability (MTSI) [48]. To demonstrate that the ion beam acts as a test beam and does not excite plasma instabilities, we consider the MTSI growth rate: the ion beam density of cm-3 results in growth rates of the order – s-1. This is much smaller than the growth rate for the lower-hybrid drift instability and therefore we discount the possibility of turbulence being driven by the ion beam.
The observation of bulk energization cannot be explained by the broadening process outlined above alone, and suggests that coherent fields which facilitate particle acceleration are generated in the plasma. A tantalizing possibility consistent with our experiment is the role of short scale electrostatic turbulence (below the interferometry diagnostic’s resolution) in developing coherent structures within the plasma. In literature, both Langmuir wave [49] and LHW [50, 51] turbulence have been shown to drive the evolution of these wave structures. On the other hand, we can also consider electromagnetic turbulence, which can be driven by plasma instabilities such as the ion Weibel instability and ion filamentation instability (IFI) [52]. To investigate this possibility, 1D particle-in-cell (PIC) simulations using the OSIRIS PIC code were performed [24], showing that the counter-streaming phase is short-lived ( ns) compared to the e-folding times of the IFI and Weibel instabilities ( ns). It is therefore unlikely that these instabilities contribute to field amplification and ion acceleration.
@@
Acknowledgements
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement no. 871124 Laserlab Europe, the UK Research and Innovation (UKRI) Frontier Research Guarantee under Grant no. EP/Y035038/1 (JETLAB), and by the Central Laser Facility Vulcan dark period community support programme 24-1.. The results presented here are based on the experiment P-22-00089, which was performed at the target station Z6 at the GSI Helmholtzzentrum fuer Schwerionenforschung, Darmstadt (Germany) in the frame of FAIR Phase-0.
Author Contributions
This project was conceived by G.G., P.T. and A.F.A.B. The experiment was designed by J.T.Y.C., J.W.D.H., G.G., P.T. and A.F.A.B. and carried out by J.T.Y.C., J.W.D.H., C.H., K.M., A.B., D.S., M.M. and H.N. The data analysis was carried out by J.T.Y.C. with support from J.W.D.H., R.B., G.G., C.H. and A.F.A.B. The manuscript was written by J.T.Y.C., with input from J.W.D.H., R.B., G.G., C.A.J.P., B.R., K.A.B. and A.F.A.B. Numerical simulations were performed by K.M. and P.T. Further experimental and theoretical support was provided by C.D.A., A.R.B., T.C., E.H., D.Q.L., F.M., P.N., A.R., S.S., A.S., C.S., C.B.S. and H.W.
Competing Interests
The authors declare no competing interests.